Pearl Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $15,000,000 on January 1, 2020. Pearl expected to complete the building by December 31, 2020. Pearl has the following debt obligations outstanding during the construction period. Construction loan-12% interest, payable semiannually, issued December 31, 2019 $6,000,000 Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2021 4,200,000 Long-term loan-11% interest, payable on January 1 of each year. Principal payable on January 1, 2024 3,000,000

Assume that Pearl completed the office and warehouse building on December 31, 2020, as planned at a total cost of $15,600,000, and the weighted-average amount of accumulated expenditures was $10,800,000. Compute the avoidable interest on this project. (Use interest rates rounded to 2 decimal places, e.g. 7.58% for computational purposes and round final answers to 0 decimal places, e.g. 5,275.)

Avoidable Interest is ?

Compute the depreciation expense for the year ended December 31, 2021. Pearl elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $900,000. (Round answer to 0 decimal places, e.g. 5,275.)

Depreciation Expense is?
Answer #1

Solution Weighted average Interest rate on general borrowinge = 10% x 420000017200000 + 11% X 3000000/7200000 = 62.92% Avoida2) Total cost of building = 15600000 + 10220160 = 75820160 Depreciation expense for 2021 = [cost – Salvage valwe] /useful lej